# Time-Dependent Black Hole Lensing and the Unified Weak-to-Strong Deflection Framework

### v1

@@ VI Conclusion
We have developed a unified, fully analytical framework that links weak-field lensing to the time-dependent strong-deflection limit for light propagating in a perturbed Schwarzschild spacetime. The backbone of our construction is the line-of-sight Born map in Eq. (24), derived from the first-order response of null geodesics to an axisymmetric even-parity $(\ell,m)=(2,0)$ QNM perturbation. By evaluating the same kernel on background rays we obtained, on one hand, the far-field expansion of Eq. (30), which produces a ringdown-frequency centroid wobble on the observer’s screen, and, on the other hand, the near-photon-sphere expansion of Eq. (38), which yields a time-dependent generalization of the logarithmic strong-deflection law. The linearized form in Eq. (41) exposes the dominance of the critical-scale modulation as $b\to b_{c}^{+}$. This single-kernel perspective ensures that weak-field and strong-field observables are phase-locked to the same QNM frequency and are related by controlled asymptotics.
@@
The imaging layer built on the observer tetrad of Eq. (9) renders all diagnostics gauge-safe at $\mathcal{O}(\varepsilon)$. In the weak field, the centroid $\bm{\theta}_{c}(t_{o})$ follows Eq. (32) and provides a clean harmonic signature. In the strong field, the hierarchy of relativistic images $\theta_{n}(t_{o})$ in Eq. (43) and the photon-ring radius $\theta_{\mathrm{ring}}(t_{o})$ in Eq. (49) exhibit coherent, frequency-resolved modulations. The logarithmic spacing $q(t_{o})$ in Eq. (50) and the inter-image delays in Eq. (48) trace the same near-photon-sphere kernel that controls the modulation of $b_{c}(t_{o})$ via Eq. (39). The matching procedure of Sec. V.1, summarized by Eq. (55), makes the continuity across regimes explicit: the residue that governs the photon-ring modulation and the coefficient that governs spacing variations are encoded in the near-critical expansion of the very $1/b$ amplitude that drives the centroid wobble. For a global, quasi-normal oscillation on a black-hole background, the leading Born imprint scales as $1/b$; for localized radiative sources, near-zone terms dominate and yield net $b^{-3}$ scaling after intermediate-wave zone cancellations.
@@
Several extensions suggest themselves while preserving the purely theoretical character of this work. First, one may include additional even-parity multipoles or odd-parity counterparts within the same Born kernel; the algebra expands, but the logic remains the same as in Eqs. (24), (30), and (38) remains unchanged. Second, although we have focused on Schwarzschild, the strategy is designed to be ported to slowly rotating backgrounds by treating spin as a controlled perturbation; the essential step is to replace the Zerilli sector with its Teukolsky analogue while keeping the observer-plane definitions fixed. For spin, the same kernel applies once the metric is reconstructed from Teukolsky variables (e.g., via the Chrzanowski-Cohen-Kegeles or Hertz potential approaches), and the observer-tetrad projection carries through unchanged. Third, the present construction invites a systematic study of phase relations among weak-field wobble, spacing, and ring modulations, such as $\Phi_{\rm c-ring}$ in Eq. (58), as robust, kernel-level diagnostics independent of source morphology. Finally, our appendices provide the algebraic infrastructure for reconstructing $h_{\mu\nu}$ from gauge-invariant master fields and for evaluating the Born kernel in both asymptotic regimes; these derivations can be adapted to explore higher-order effects without resorting to numerical implementations.
@@
To end, by anchoring all lensing observables to a single time-dependent Born representation and by tracing their weak-to-strong continuity, we have established an analytic, gauge-controlled framework for QNM-driven imaging phenomenology around Schwarzschild black holes. The resulting relations among centroid motion, multi-image structure, and photon-ring modulations supply a compact theoretical toolbox that is readily extensible and that clarifies how ringdown physics imprints across the entire screen.
@@ Acknowledgements.
A. O and R. P. would like to acknowledge networking support of the COST Action CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers), the COST Action CA22113 - Fundamental challenges in theoretical physics (THEORY-CHALLENGES), the COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), the COST Action CA23130 - Bridging high and low energies in search of quantum gravity (BridgeQG), and the COST Action CA23115 - Relativistic Quantum Information (RQI) funded by COST (European Cooperation in Science and Technology). A. O and R. P. would also like to acknowledge the funding support of SCOAP3.
@@ Appendix A Zerilli framework and reconstruction to $h_{\mu\nu}$
In this appendix, we collect the even-parity RWZ machinery used in Secs. II-V. We keep the background Schwarzschild metric $g^{(0)}_{\mu\nu}$ and notation of Eq. (1), the perturbative split of Eq. (3), and the Zerilli master equation and potential of Eqs. (6)-(7), together with the QNM boundary conditions in Eq. (8). Our goal is twofold: to define a gauge-invariant even-parity master field $\Psi_{Z}$ appropriate for $(\ell,m)=(2,0)$, and to reconstruct the RWZ-gauge metric amplitudes $H_{0},H_{1},H_{2},K$ and the combination $h_{\alpha\beta}p^{\alpha}p^{\beta}$ that appears in the Born kernel of Eqs. (20)-(24).
@@
$h_{AB}$ $=r^{2}K^{\ell m}(t,r)\,\Omega_{AB}\,Y_{\ell m},$ (65)
with $A,B\in{\theta,\phi}$, $\Omega_{AB}=\mathrm{diag}(1,\sin^{2}\theta)$, and f=12M/rf=1-2M/r as in Eq. (1). The RWZ gauge conditions set the even-parity vector and tensor harmonics $(j_{a},G)$ to zero. For axisymmetry ($m=0$) we have $Y_{\ell 0}(\theta)=\sqrt{\frac{2\ell+1}{4\pi}}\,P_{\ell}(\cos\theta)$. All amplitudes $H_{0},H_{1},H_{2},K$ used in the main text (Secs. II-V) are defined in the $P_{\ell}$ basis. When needed, $Y$-basis amplitudes and $P$-basis amplitudes are related by $H_{a}^{(Y)}=\sqrt{\tfrac{2\ell+1}{4\pi}}\,H_{a}^{(P)}$ for $a\in\{0,1,2,K\}$. It is also convenient to adopt the tortoise coordinate $r_{*}$ of Eq. (5), so that the wave operator acting on master fields takes the flat-space form in $(t,r_{*})$.
@@
$\lambda$ $\equiv\frac{(\ell-1)(\ell+2)}{2}=2.$ (66)
By construction, $\Psi_{Z}$ is invariant under linearized even-parity gauge transformations [cf. Eq. (15)]. In vacuum it satisfies the Zerilli equation already quoted in Eq. (6),
$\left(-\partial_{t}^{2}+\partial_{r_{*}}^{2}-V_{Z}(r)\right)\Psi_{Z}=0,$ (67)
with $V_{Z}$ given in Eq. (7). For QNMs, we impose the homogeneous boundary conditions of Eq. (8). In the frequency domain, $\Psi_{Z}(t,r)=\psi(r)\,e^{-i\omega t}$ with complex $\omega=\omega_{R}+i\omega_{I}$, $\omega_{I}<0$.
A normalization convention is needed to separate the physical smallness of the perturbation from the shape of $\Psi_{Z}$. Throughout we take the outgoing amplitude of $\Psi_{Z}$ at large $r$ to be unity and collect the overall smallness in the book-keeping parameter $\varepsilon$ of Eq. (3); this is the convention used in Secs. III-V.
Reconstruction of $H_{0},H_{1},H_{2},K$ from $\Psi_{Z}$: In vacuum, the Einstein equations relate $(H_{0},H_{1},H_{2},K)$ algebraically to $\Psi_{Z}$ and its first derivatives. One convenient RWZ-gauge reconstruction, obtained by inverting Eq. (66) and using the field equations, reads
@@
$H_{1}$ $=-\frac{i\omega\,\mathcal{N}_{1}(r)}{f(\lambda r+3M)}\,\Psi_{Z}+i\omega\,r,\partial_{r}\Psi_{Z},$ (68)
@@
$\mathcal{N}_{1}$ $=\lambda r^{2}+3\lambda Mr+6M^{2}.$ (69)
@@
We remark that (i) The equality $H_{0}=H_{2}$ follows from the vacuum field equations in RWZ gauge; (ii) Other algebraically equivalent reconstructions exist; any such choice leads to the same gauge-invariant observables; (iii) The particular coefficients in Eq. (69) are tailored to the definition in Eq. (66). If a different master variable is adopted, the polynomials change accordingly without affecting physical content.
Asymptotics and horizon behavior: From Eq. (8), $\Psi_{Z}$ behaves as an outgoing wave at infinity and as an ingoing wave at the horizon,
$\Psi_{Z}\sim\begin{cases}A_{\infty}\,e^{-i\omega(t-r_{*})},&r\to\infty,\\ A_{H}\,e^{-i\omega(t+r_{*})},&r\to 2M,\end{cases}$ (70)
with $A_{\infty}=1$ by our convention and $A_{H}$ determined by the solution of Eq. (67). Substituting $A.6$ into $A.4$ and using $r_{*}\sim r+2M\ln(r/2M-1)$, we find
@@
$H_{0,1,2},K$ $=\mathcal{O}(1)\,e^{-i\omega(t+r_{*})}\quad\text{for}\quad(r\to 2M),$ (71)
@@
The contraction $h_{\alpha\beta}p^{\alpha}p^{\beta}$ along a background ray: The Born kernel in Eqs. (20)-(24) involves $\partial_{\mu}h_{\alpha\beta}\,p_{0}^{\alpha}p_{0}^{\beta}$ evaluated on the Schwarzschild null geodesic $x_{0}^{\mu}(\lambda;b)$. Using Eq. (65) with $m=0$ and the axisymmetric harmonic $Y_{20}(\theta)=\sqrt{5/(4\pi)}\,P_{2}(\cos\theta)$, we obtain
@@
$+f^{-1}H_{2}\,(p_{0}^{r})^{2}+r^{2}K\,\Omega_{AB},\hat{p}_{0}^{A}\hat{p}_{0}^{B}\Bigr],$ (72)
where $\hat{p}_{0}^{A}=(p_{0}^{\theta}\,p_{0}^{\phi})$ are the angular components of the background momentum and $\Omega_{AB}\,\hat{p}_{0}^{A}\hat{p}_{0}^{B}=(p_{0}^{\theta})^{2}+\sin^{2}\theta,(p_{0}^{\phi})^{2}=L^{2}/r^{4}$ with $L$ the conserved angular momentum of the background geodesic [cf. Eq. (14)]. Substituting Eq. (68) into Eq. (72) and simplifying gives the compact representation
$h_{\alpha\beta}p_{0}^{\alpha}p_{0}^{\beta}=Y_{20}(\theta)\left[\mathcal{P}_{0}(r;E,L)\,\Psi_{Z}+\mathcal{P}_{1}(r;E,L)\,\partial_{r}\Psi_{Z}\right],$ (73)
@@
$=\frac{f}{r}\,(E^{2}-fL^{2}/r^{2})+2i\omega r\,E\,\sqrt{E^{2}-fL^{2}/r^{2}},$ (74)
where $\mathcal{A}_{0}(r)\equiv\mathcal{N}_{K}(r)+\mathcal{N}_{2}(r)$ and we have expressed the radial momentum as $p_{0}^{r}=\pm\sqrt{E^{2}-fL^{2}/r^{2}}$ with $E=-p_{0t}$ the conserved energy (the sign flips at turning points). Eqs. (73)-(74) make explicit that only $\Psi_{Z}$ and $\partial_{r}\Psi_{Z}$ enter the kernel, and that the dependence on the orbital constants $(E,L)$ factorizes algebraically.
Taking a derivative and using the background geodesic equations, the source term in Eq. (18) can be written in the equivalent forms
@@
$=\left(-i\omega+\frac{dr_{*}}{d\lambda}\,\partial_{r_{*}}\right)\left[h_{\alpha\beta}p_{0}^{\alpha}p_{0}^{\beta}\right]\,d\lambda\,$ (75)
which is useful for casting the line-of-sight integrals (20)-(24) in terms of the tortoise coordinate $r_{*}$ and the delay $\tau$ used in Eq. (24).
Large-(r) limit and normalization check: With the outgoing normalization $A_{\infty}=1$ in Eq. (70), the asymptotic behavior of Eq. (68) implies
@@
$\times\left[\mathfrak{c}_{0}(\omega;E,L)+\mathfrak{c}_{1}(\omega;E,L)\,\mathcal{O}(1/r)\right],$ (76)
for some algebraic coefficients $\mathfrak{c}_{0,1}$. Substitution of Eq. (76) into the kernel Eq. (21) reproduces the far-zone scaling used in Sec. III.2 and leads directly to the $1/b$ dependence in Eq. (27), providing a consistency check of the normalization choices above.
Near-photon-sphere reduction: Near the photon sphere $r_{\mathrm{ph}}=3M$ [Eq. (33)], the master equation in Eq. (67) can be approximated by a Schrödinger problem with an inverted harmonic potential around the unstable maximum of $V_{Z}$. Expanding $\Psi_{Z}$ and its derivative in that region and inserting into Eq. (68) and then (A.8) yields
@@
$+\mathfrak{C}_{1}(\omega;E,L)\,\partial_{r}\Psi_{Z}\Bigr]_{r=3M}+\ldots,$ (77)
with constants $\mathfrak{C}_{0,1}$ determined by the polynomials in Eq. (69). This reduction underlies the logarithmic structures extracted in Sec. IV, culminating in Eqs. (37)-(41).
Eqs. (66)-(69) provide a self-contained RWZ-gauge reconstruction from the Zerilli-Moncrief master field to the metric amplitudes $(H_{0},H_{1},H_{2},K)$. The contraction formulas in Eqs. (72)-(75) deliver the precise input required by the Born kernel of Eq. (21), and their far-zone and near-photon-sphere limits, Eqs. (76)-(77), are the only ingredients needed to derive the weak-field expansion (Sec. III.2) and the time-dependent SDL laws (Sec. IV). All steps are consistent with the QNM boundary conditions in Eq. (8), and all observable statements inherit gauge safety from the use of $\Psi_{Z}$ and the observer tetrad defined in Eq. (9).

### v2

@@ VI Conclusion
We have constructed an analytic description that connects weak-field lensing to a time-dependent strong-deflection limit for light propagating in a perturbed Schwarzschild spacetime. The backbone of our construction is the line-of-sight Born map in Eq. (III.7), derived from the first-order response of null geodesics to an axisymmetric even-parity $(\ell,m)=(2,0)$ QNM perturbation. By evaluating the same kernel on background rays we obtained, on one hand, the far-field expansion of Eq. (III.13), which produces a ringdown-frequency centroid wobble on the observer’s screen, and, on the other hand, the near-photon-sphere expansion of Eq. (IV.6), which yields a time-dependent generalization of the logarithmic strong-deflection law. The linearized form in Eq. (IV.9) exposes the dominance of the critical-scale modulation as $b\to b_{c}^{+}$. This single-kernel perspective makes explicit that weak-field and strong-field observables are phase-locked to the same QNM frequency and are related by controlled asymptotics.
@@
The imaging layer built on the observer tetrad of Eq. (II.9) renders all diagnostics gauge-safe at $\mathcal{O}(\varepsilon)$. In the weak field, the centroid $\bm{\theta}_{c}(t_{o})$ follows Eq. (III.15) and provides a clean harmonic signature. In the strong field, the hierarchy of relativistic images $\theta_{n}(t_{o})$ in Eq. (IV.13) and the photon-ring radius $\theta_{\mathrm{ring}}(t_{o})$ in Eq. (IV.19) exhibit coherent, frequency-resolved modulations. The logarithmic spacing $q(t_{o})$ in Eq. (IV.20) and the inter-image delays in Eq. (IV.18) trace the same near-photon-sphere kernel that controls the modulation of $b_{c}(t_{o})$ via Eq. (IV.7). The matching procedure of Sec. V.1, summarized by Eq. (V.5), makes the continuity across regimes explicit: the residue that governs the photon-ring modulation and the coefficient that governs spacing variations are encoded in the near-critical expansion of the very $1/b$ amplitude that drives the centroid wobble. For a global, quasi-normal oscillation on a black-hole background, the leading Born imprint scales as $1/b$; for localized radiative sources, near-zone terms dominate and yield net $b^{-3}$ scaling after intermediate-wave zone cancellations.
@@
Several extensions can be pursued within the same analytic setup. One may include additional even-parity multipoles, or odd-parity counterparts, by incorporating the corresponding gauge-invariant master fields into the Born kernel; the derivations follow the same steps that lead to Eqs. (III.7), (III.13), and (IV.6). Although we have focused on Schwarzschild, the strategy can be generalized to slowly rotating backgrounds by treating the spin as a controlled perturbation, replacing the Zerilli sector with its Teukolsky analogue while keeping the observer-plane definitions fixed. The present construction also motivates a systematic study of phase relations among weak-field wobble, spacing, and ring modulations, such as $\Phi_{\rm c-ring}$ in Eq. (V.8), as kernel-level diagnostics that are insensitive to source morphology. Finally, the appendices collect the reconstruction and kernel-evaluation steps used here and can be adapted to higher-order corrections when these remain under analytic control.
@@
To end, expressing the deflection and the associated imaging observables in terms of a single time-dependent Born representation makes the weak-to-strong continuity transparent and keeps the gauge control explicit. The resulting relations among centroid motion, multi-image structure, and photon-ring modulations clarify how ringdown-scale perturbations enter screen observables around Schwarzschild black holes.
@@ Acknowledgements.
The author is grateful for helpful discussions with A. Övgün. R. P. would like to acknowledge networking support of the COST Action CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers), the COST Action CA22113 - Fundamental challenges in theoretical physics (THEORY-CHALLENGES), the COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), the COST Action CA23130 - Bridging high and low energies in search of quantum gravity (BridgeQG), the COST Action CA23115 - Relativistic Quantum Information (RQI) funded by COST (European Cooperation in Science and Technology), and SCOAP3 (Switzerland) for their support.
@@ Appendix A Zerilli framework and reconstruction to $h_{\mu\nu}$
In this appendix, we collect the even-parity RWZ machinery used in Secs. II-V. We keep the background Schwarzschild metric $g^{(0)}_{\mu\nu}$ and notation of Eq. (II.1), the perturbative split of Eq. (II.3), and the Zerilli master equation and potential of Eqs. (II.6)-(II.7), together with the QNM boundary conditions in Eq. (II.8). Our goal is twofold: to define a gauge-invariant even-parity master field $\Psi_{Z}$ appropriate for $(\ell,m)=(2,0)$, and to reconstruct the RWZ-gauge metric amplitudes $H_{0},H_{1},H_{2},K$ and the combination $h_{\alpha\beta}p^{\alpha}p^{\beta}$ that appears in the Born kernel of Eqs. (III.4)-(III.7).
@@
$h_{AB}$ $=r^{2}K^{\ell m}(t,r)\,\Omega_{AB}\,Y_{\ell m},$ (A.1)
with $A,B\in{\theta,\phi}$, $\Omega_{AB}=\mathrm{diag}(1,\sin^{2}\theta)$, and f=12M/rf=1-2M/r as in Eq. (II.1). The RWZ gauge conditions set the even-parity vector and tensor harmonics $(j_{a},G)$ to zero. For axisymmetry ($m=0$) we have $Y_{\ell 0}(\theta)=\sqrt{\frac{2\ell+1}{4\pi}}\,P_{\ell}(\cos\theta)$. All amplitudes $H_{0},H_{1},H_{2},K$ used in the main text (Secs. II-V) are defined in the $P_{\ell}$ basis. When needed, $Y$-basis amplitudes and $P$-basis amplitudes are related by $H_{a}^{(Y)}=\sqrt{\tfrac{2\ell+1}{4\pi}}\,H_{a}^{(P)}$ for $a\in\{0,1,2,K\}$. It is also convenient to adopt the tortoise coordinate $r_{*}$ of Eq. (II.5), so that the wave operator acting on master fields takes the flat-space form in $(t,r_{*})$.
@@
$\lambda$ $\equiv\frac{(\ell-1)(\ell+2)}{2}=2.$ (A.2)
By construction, $\Psi_{Z}$ is invariant under linearized even-parity gauge transformations [cf. Eq. (II.15)]. In vacuum it satisfies the Zerilli equation already quoted in Eq. (II.6),
$\left(-\partial_{t}^{2}+\partial_{r_{*}}^{2}-V_{Z}(r)\right)\Psi_{Z}=0,$ (A.3)
with $V_{Z}$ given in Eq. (II.7). For QNMs, we impose the homogeneous boundary conditions of Eq. (II.8). In the frequency domain, $\Psi_{Z}(t,r)=\psi(r)\,e^{-i\omega t}$ with complex $\omega=\omega_{R}+i\omega_{I}$, $\omega_{I}<0$.
A normalization convention is needed to separate the physical smallness of the perturbation from the shape of $\Psi_{Z}$. Throughout we take the outgoing amplitude of $\Psi_{Z}$ at large $r$ to be unity and collect the overall smallness in the book-keeping parameter $\varepsilon$ of Eq. (II.3); this is the convention used in Secs. III-V.
Reconstruction of $H_{0},H_{1},H_{2},K$ from $\Psi_{Z}$: In vacuum, the Einstein equations relate $(H_{0},H_{1},H_{2},K)$ algebraically to $\Psi_{Z}$ and its first derivatives. One convenient RWZ-gauge reconstruction, obtained by inverting Eq. (A.2) and using the field equations, reads
@@
$H_{1}$ $=-\frac{i\omega\,\mathcal{N}_{1}(r)}{f(\lambda r+3M)}\,\Psi_{Z}+i\omega\,r,\partial_{r}\Psi_{Z},$ (A.4)
@@
$\mathcal{N}_{1}$ $=\lambda r^{2}+3\lambda Mr+6M^{2}.$ (A.5)
@@
We remark that (i) The equality $H_{0}=H_{2}$ follows from the vacuum field equations in RWZ gauge; (ii) Other algebraically equivalent reconstructions exist; any such choice leads to the same gauge-invariant observables; (iii) The particular coefficients in Eq. (A.5) are tailored to the definition in Eq. (A.2). If a different master variable is adopted, the polynomials change accordingly without affecting physical content.
Asymptotics and horizon behavior: From Eq. (II.8), $\Psi_{Z}$ behaves as an outgoing wave at infinity and as an ingoing wave at the horizon,
$\Psi_{Z}\sim\begin{cases}A_{\infty}\,e^{-i\omega(t-r_{*})},&r\to\infty,\\ A_{H}\,e^{-i\omega(t+r_{*})},&r\to 2M,\end{cases}$ (A.6)
with $A_{\infty}=1$ by our convention and $A_{H}$ determined by the solution of Eq. (A.3). Substituting $A.6$ into $A.4$ and using $r_{*}\sim r+2M\ln(r/2M-1)$, we find
@@
$H_{0,1,2},K$ $=\mathcal{O}(1)\,e^{-i\omega(t+r_{*})}\quad\text{for}\quad(r\to 2M),$ (A.7)
@@
The contraction $h_{\alpha\beta}p^{\alpha}p^{\beta}$ along a background ray: The Born kernel in Eqs. (III.4)-(III.7) involves $\partial_{\mu}h_{\alpha\beta}\,p_{0}^{\alpha}p_{0}^{\beta}$ evaluated on the Schwarzschild null geodesic $x_{0}^{\mu}(\lambda;b)$. Using Eq. (A.1) with $m=0$ and the axisymmetric harmonic $Y_{20}(\theta)=\sqrt{5/(4\pi)}\,P_{2}(\cos\theta)$, we obtain
@@
$+f^{-1}H_{2}\,(p_{0}^{r})^{2}+r^{2}K\,\Omega_{AB},\hat{p}_{0}^{A}\hat{p}_{0}^{B}\Bigr],$ (A.8)
where $\hat{p}_{0}^{A}=(p_{0}^{\theta}\,p_{0}^{\phi})$ are the angular components of the background momentum and $\Omega_{AB}\,\hat{p}_{0}^{A}\hat{p}_{0}^{B}=(p_{0}^{\theta})^{2}+\sin^{2}\theta,(p_{0}^{\phi})^{2}=L^{2}/r^{4}$ with $L$ the conserved angular momentum of the background geodesic [cf. Eq. (II.14)]. Substituting Eq. (A.4) into Eq. (A.8) and simplifying gives the compact representation
$h_{\alpha\beta}p_{0}^{\alpha}p_{0}^{\beta}=Y_{20}(\theta)\left[\mathcal{P}_{0}(r;E,L)\,\Psi_{Z}+\mathcal{P}_{1}(r;E,L)\,\partial_{r}\Psi_{Z}\right],$ (A.9)
@@
$=\frac{f}{r}\,(E^{2}-fL^{2}/r^{2})+2i\omega r\,E\,\sqrt{E^{2}-fL^{2}/r^{2}},$ (A.10)
where $\mathcal{A}_{0}(r)\equiv\mathcal{N}_{K}(r)+\mathcal{N}_{2}(r)$ and we have expressed the radial momentum as $p_{0}^{r}=\pm\sqrt{E^{2}-fL^{2}/r^{2}}$ with $E=-p_{0t}$ the conserved energy (the sign flips at turning points). Eqs. (A.9)-(A.10) make explicit that only $\Psi_{Z}$ and $\partial_{r}\Psi_{Z}$ enter the kernel, and that the dependence on the orbital constants $(E,L)$ factorizes algebraically.
Taking a derivative and using the background geodesic equations, the source term in Eq. (III.2) can be written in the equivalent forms
@@
$=\left(-i\omega+\frac{dr_{*}}{d\lambda}\,\partial_{r_{*}}\right)\left[h_{\alpha\beta}p_{0}^{\alpha}p_{0}^{\beta}\right]\,d\lambda\,$ (A.11)
which is useful for casting the line-of-sight integrals (III.4)-(III.7) in terms of the tortoise coordinate $r_{*}$ and the delay $\tau$ used in Eq. (III.7).
Large-(r) limit and normalization check: With the outgoing normalization $A_{\infty}=1$ in Eq. (A.6), the asymptotic behavior of Eq. (A.4) implies
@@
$\times\left[\mathfrak{c}_{0}(\omega;E,L)+\mathfrak{c}_{1}(\omega;E,L)\,\mathcal{O}(1/r)\right],$ (A.12)
for some algebraic coefficients $\mathfrak{c}_{0,1}$. Substitution of Eq. (A.12) into the kernel Eq. (III.5) reproduces the far-zone scaling used in Sec. III.2 and leads directly to the $1/b$ dependence in Eq. (III.10), providing a consistency check of the normalization choices above.
Near-photon-sphere reduction: Near the photon sphere $r_{\mathrm{ph}}=3M$ [Eq. (IV.1)], the master equation in Eq. (A.3) can be approximated by a Schrödinger problem with an inverted harmonic potential around the unstable maximum of $V_{Z}$. Expanding $\Psi_{Z}$ and its derivative in that region and inserting into Eq. (A.4) and then (A.8) yields
@@
$+\mathfrak{C}_{1}(\omega;E,L)\,\partial_{r}\Psi_{Z}\Bigr]_{r=3M}+\ldots,$ (A.13)
with constants $\mathfrak{C}_{0,1}$ determined by the polynomials in Eq. (A.5). This reduction underlies the logarithmic structures extracted in Sec. IV, culminating in Eqs. (IV.5)-(IV.9).
Eqs. (A.2)-(A.5) provide a self-contained RWZ-gauge reconstruction from the Zerilli-Moncrief master field to the metric amplitudes $(H_{0},H_{1},H_{2},K)$. The contraction formulas in Eqs. (A.8)-(A.11) deliver the precise input required by the Born kernel of Eq. (III.5), and their far-zone and near-photon-sphere limits, Eqs. (A.12)-(A.13), are the only ingredients needed to derive the weak-field expansion (Sec. III.2) and the time-dependent SDL laws (Sec. IV). All steps are consistent with the QNM boundary conditions in Eq. (II.8), and all observable statements inherit gauge safety from the use of $\Psi_{Z}$ and the observer tetrad defined in Eq. (II.9).